Two objects moving along same straight line are leaving point A with an acceleration a, 2a & velocity 2u, u respectively at time t=0. The distance moved by the object with respect to point A when one object overtake the other is

Grade:11

1 Answers

Arun
6 years ago
Dear Student
 
 

When one object will overtake the other, both of them will travel the same distance in the same time interval.

 

Distance travelled by the first body in a time interval ''t''(say) ,will be, by the scond kinematic equation,

 

d = 2ut+1/2at^2

 

Also, distance travelled by the second body will be

 

d=ut+1/2*2at^2

 

Equating the two, we get

 

2ut+1/2at^2=ut+at^2

 

Solving, we get t=0 and t=2u/a

 

d=2u*2u/a+1/2a*4u^2/a^2

 

d=6u^2/a

 

Regards

Arun (askIITians forum expert)